The carpenter saws a board into two pieces. He wants one piece to be 3 feet longer than twice the length of the shorter piece. Find the length of each piece when a = 30.

Answer by mananth(16946)

The carpenter saws a board into two pieces. He wants one piece to be 3 feet longer than twice the length of the shorter piece. Find the length of each piece when a = 30.
...
let the shorter piece be x
the larger piece = 2x+3
x+2x+3=30
3x+3=30
-3
3x=27
x= 9
..
one piece = 9 other 21 feet
..
m.ananth@hotmail.ca